• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fatiha

    Log entry time 08:02:23 on March 1,2009

    Entry number 264013

    keyword=Shift Summary

    Shift crew: F. Benmokhtar, B. Michaels and R. Lindgren
    ----------

    00:42 Beam down, RF cavity issues, MCC is calling in the
    Techniciens. It will take a coupe of hours to restore
    the beam.
    2:00 Brad and Kalyan went for a controlled access to work
    on the trigger.
    3:10 MCC called, it lookes like they've solved the problems
    with the RF cavities but having some issues with the
    liquid level in the cryos. They called and expert.
    3:27 MCC called, says they are done with what they wanted to
    do.
    3:35 MCC doign harp scan.
    4:05 Moved to Be) target, ask for raster on, 2*2.6 (MCC units)
    2uA.
    Got good beam positions, spot++ 2 by 2.
    4:14 3He, 2uA. rates consistant between 2*2.6 and 5.2*6.2,
    Note that MCC couldn't set 5.2*6.5 (reaches the limit)
    4:30 Back to production running, Kin = I, 5:05 Bigbite S14, channels 14.6 and 14.7 HV trip. Recovered.
    run# 1912
    5:40 Beam down, beam pipe vaccum problem.
    5:45 We asked for a short controlled access. Brad and Kalyan
    going down for setting the trigger timing.
    6:03 Back from controlled access asked for beam back.

    Bigbite Run list:
    ----------------
    1909 -- 3He Ended, beam down
    1911 -- // Ended, beam pipe problem
    1912 -- // Prescale test
    1913 -- // Ended, beam pipe vaccum issues.
    1914 -- First complete production run.

    LHRS run list:
    --------------
    20461 -- Ended, beam down very short run
    20462 -- BeO target, beam position cehck
    20463 -- 3He, short run, ended beam pipe issues.
    20464 -- 3He, production
    20465 -- 3He, Ended, beam pipe vaccum issues.
    20466 -- First complete production run.